OpenStudy (anonymous):

What's the derivative of the function y=f(2+lnx) ? The options are: A) y'=2/x*f'(2+lnx) B) y'=1/x C) y'=1/x*f'(2+lnx) D) y'=2+lnx

OpenStudy (anonymous):

Differentiate both sides, but be careful to use the chain rule correctly on the right hand side. This will give you:\[y'=\frac{df}{dx}\frac{d(2+lnx)}{dx}\]

OpenStudy (anonymous):

Which is simply \[y'=\frac{1}{x}f'\], or choice C

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

Sure thing